OpenStudy (anonymous):

Which statement is true about the product square root of 2(5 + square root of 8)? It is rational and equal to 7. It is rational and equal to 9. It is irrational and equal to 4 + square root of 10. It is irrational and equal to 4 + 5square root of 2.

OpenStudy (anonymous):

\[\sqrt{2}(5+\sqrt{8}\]

OpenStudy (anonymous):

forgot the last parenth

OpenStudy (tkhunny):

Hint: \(\sqrt{8} = 2\sqrt{2}\)

OpenStudy (anonymous):

this wouldnt be right? \[5\sqrt{2}+\sqrt{16}\] \[4+5\sqrt{2}\]

OpenStudy (anonymous):

@tkhunny ?

OpenStudy (tkhunny):

You didn't use my hint. Try to keep things simpler. you don't have to face that 16 if you don't want to. In any case, you have it right. Now, answer the question.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

ty for the help :) sorry for not using your hint i tried it and got really confused then tried with the 16 and saw it matched with one of the answers

OpenStudy (tkhunny):

No worries. I have no such austerity. You finding a correct solution and understanding what you are doing is my most important concern.